# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@ 4.6. Null geodesics of $g^{\mathrm{F}}$
Since $\widehat{f}$ maps any (null) chain on $M_{1}$ to a (null) chain on $M_{2}$, $f$ maps conformal geodesics on $\Sigma_{1}$ to conformal geodesics on $\Sigma_{2}$. Hence, by [16, Theorem 2.3], $f$ is a conformal isomorphism, ∎
4.7. Canonical lift of conformal geodesics
Theorem 1.1 (2) asserts that we can lift a conformal geodesic $x(t)$ to a unique (null) chain on $M$ once we choose a lift of the endpoint $x(a)$; we prove this in §5.1 below. Here, as a special case, we present a canonical lift of a conformal geodesic to a chain.
Given a regular curve $x(t)$ on $\Sigma$, we take a complex vector field $\zeta(t)$ along $x(t)$ such that $(\sqrt{2}\,\mathrm{Re}\,\zeta,\sqrt{2}\,\mathrm{Im}\,\zeta)$ gives an oriented orthonormal basis of $\dot{x}^{\perp}$ with respect to a metric $g\in[g]$. Such a $\zeta(t)$ is determined by the conformal structure $[g]$ up to $\mathbb{C}^{*}$-action, so we have a canonical lift
@@ 4.7. Canonical lift of conformal geodesics
The curve $\tilde{x}(t)$ is a chain if and only if $x(t)$ is a conformal geodesic. Moreover, for a conformal geodesic $x(t)\ (a\leq t\leq b)$, the curve $\tilde{x}(t)$ is a unique chain through $(x(a),[\zeta(a)])$ which projects to $x(t)$.
@@
$\eta=\dot{x},\quad\dot{\gamma}^{0}=1,\quad\dot{\gamma}^{1}=\dot{\gamma}^{2}=\dot{\gamma}^{3}=0.$
@@
Next, we prove the uniqueness. Suppose that $c(t)$ is a chain through $(x(a),[\zeta(a)])$ which projects to a conformal geodesic $x(t)$, and let $\gamma_{c}(t)=(x(t),\zeta_{c}(t))$ be a null geodesic on $(\widehat{\mathcal{C}},g^{\mathrm{F}})$, the Fefferman space determined by $g$ above, which projects to $c(t)$. Then, $\dot{\gamma}_{c}^{0}=g(\eta_{c},\dot{x})$ and $\dot{\gamma}_{c}^{1}=g(\zeta_{c},\dot{x})$ are constant by the geodesic equation. Since $\dot{\gamma}_{c}^{0}(a)=|\dot{x}(a)|^{2}=1$, $\dot{\gamma}_{c}^{1}(a)=0$, we obtain that $\eta_{c}(t)=\dot{x}(t)$ for any $t$ and hence $c(t)=\tilde{x}(t)$. ∎
5. A variational principle for conformal geodesics in dimension three
5.1. Relations to the unit tangent sphere bundle and the frame bundle
We will give an alternative description of the previous constructions in terms of the frame bundle over $\Sigma$.
We fix a representative metric $g\in[g]$ and let $\mathcal{P}$ be the $\mathrm{SO}(3)$-bundle of oriented orthonormal frames over $(\Sigma,g)$. Then, the map
(5.1) $f\colon\widehat{\mathcal{C}}\overset{\sim}{\longrightarrow}\mathcal{P},\quad\zeta\longmapsto B=(\sqrt{2}\,{\rm Re}\,\zeta,\sqrt{2}\,{\rm Im}\,\zeta,\eta)$
@@ 5.1. Relations to the unit tangent sphere bundle and the frame bundle
(5.2) $M\overset{\sim}{\longrightarrow}S\Sigma,\quad[\zeta]\longmapsto\eta.$
@@
$\phi=\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix}\colon T\mathcal{P}\longrightarrow\mathbb{R}^{3},\quad\omega=\begin{pmatrix}0&\omega^{3}&-\omega^{2}\\ -\omega^{3}&0&\omega^{1}\\ \omega^{2}&-\omega^{1}&0\end{pmatrix}\colon T\mathcal{P}\longrightarrow\mathfrak{so}(3)$
be the canonical $1$-form and the Levi-Civita connection form of $g$. These forms are related to the adapted coframe $\{\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\}$ on $\widehat{\mathcal{C}}$ as follows:
@@
$f^{*}\phi=\begin{pmatrix}\sqrt{2}\,{\rm Re}\,\tilde{\theta}^{1}\\ \sqrt{2}\,{\rm Im}\,\tilde{\theta}^{1}\\ \theta\end{pmatrix},\quad f^{*}\omega=\begin{pmatrix}0&\tilde{\theta}^{3}&-\sqrt{2}\,{\rm Im}\,\tilde{\theta}^{2}\\ -\tilde{\theta}^{3}&0&\sqrt{2}\,{\rm Re}\,\tilde{\theta}^{2}\\ \sqrt{2}\,{\rm Im}\,\tilde{\theta}^{2}&-\sqrt{2}\,{\rm Re}\,\tilde{\theta}^{2}&0\end{pmatrix}.$
@@
$f^{*}\phi={}^{t}\!B\,dx=\begin{pmatrix}\sqrt{2}\,{\rm Re}\,{}^{t}\zeta\\ \sqrt{2}\,{\rm Im}\,{}^{t}\zeta\\ {}^{t}\eta\end{pmatrix}\begin{pmatrix}dx^{1}\\ dx^{2}\\ dx^{3}\end{pmatrix}$
@@
$f^{*}\omega={}^{t}\!B\,dB=\begin{pmatrix}\sqrt{2}\,{\rm Re}\,{}^{t}\zeta\\ \sqrt{2}\,{\rm Im}\,{}^{t}\zeta\\ {}^{t}\eta\end{pmatrix}\begin{pmatrix}\sqrt{2}\,{\rm Re}\,d\zeta&\sqrt{2}\,{\rm Im}\,d\zeta&d\eta\end{pmatrix},$
@@
We define the Fefferman metric $G^{\mathrm{F}}$ on $\mathcal{P}$ via the isomorphism (5.1): $g^{\mathrm{F}}=f^{*}G^{\mathrm{F}}$. Then, the above lemma and Theorem 4.5 prove the following
@@
$G^{\mathrm{F}}=2(\phi^{1}\cdot\omega^{1}+\phi^{2}\cdot\omega^{2}+\phi^{3}\cdot\omega^{3}).$
@@
(i) If we denote by $R_{A}$ the right action of $A\in\mathrm{SO}(3)$ on $\mathcal{P}$, we have
$R_{A}^{*}\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix}=A^{-1}\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix},\quad R_{A}^{*}\begin{pmatrix}\omega^{1}\\ \omega^{2}\\ \omega^{3}\end{pmatrix}=A^{-1}\begin{pmatrix}\omega^{1}\\ \omega^{2}\\ \omega^{3}\end{pmatrix}.$
@@
(ii) Since the conformal class of the Fefferman metric is invariantly defined from the CR structure, the conformal manifold $(\mathcal{P},[G^{\mathrm{F}}])$ should also be canonically associated to $(\Sigma,[g])$. This can be checked directly as follows. Let $\widehat{g}=e^{2\Upsilon}g$ be another representative metric and $\widehat{\mathcal{P}}$ be the oriented orthonormal frame bundle for $\widehat{g}$. Then, we have the canonical bundle isomorphism
$\delta\colon\mathcal{P}\overset{\sim}{\longrightarrow}\widehat{\mathcal{P}},\quad B\longmapsto e^{-\Upsilon}B.$
We denote by $\widehat{\phi},\widehat{\omega}$ the canonical 1-form and the connection form of $\widehat{g}$ on $\widehat{\mathcal{P}}$. Then, we have $\delta^{*}\widehat{\phi}=e^{\Upsilon}\phi$ and the conformal transformation formula for the Levi-Civita connection gives
@@
$\delta^{*}\widehat{G}^{\mathrm{F}}=e^{\Upsilon}G^{\mathrm{F}}.$
In §4.7, we defined a canonical lift of a regular curve $x(t)$ on $\Sigma$ to a curve $\tilde{x}(t)=(x(t),[\zeta(t)])$ on $M$ which becomes a chain if and only if $x(t)$ is a conformal geodesic. This lift corresponds to the curve $(x(t),\dot{x}(t)/|\dot{x}(t)|)$ on $S\Sigma$ via the isomorphism (5.2). Thus, if we regard $S\Sigma$ as the twistor CR manifold, Theorem 4.12 yields
@@
For a regular curve $x(t)$ on $(\Sigma,g)$, the curve $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$ on $S\Sigma$ is a chain if and only if $x(t)$ is a conformal geodesic of $[g]$. Moreover, for a conformal geodesic $x(t)\ (a\leq t\leq b)$, the curve $\tilde{x}(t)$ is a unique chain through $(x(a),\dot{x}(a)/|\dot{x}(a)|)$ which projects to $x(t)$.
We can compete the proof of Theorem 1.1 (2) by making use of the right $\mathrm{SO}(3)$-invariance of $G^{\mathrm{F}}$:
@@
Let $x(t)\ (a\leq t\leq b)$ be a conformal geodesic on $\Sigma$ and $\eta_{0}\in S_{x(a)}\Sigma$. Take a null geodesic $\gamma(t)$ on $(\mathcal{P},G^{\mathrm{F}})$ which projects to the chain $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$. We choose $A\in\mathrm{SO}(3)$ so that the third basis vector in $R_{A}(\gamma(a))$ coincides with $\eta_{0}$, and set $\gamma^{\prime}(t):=R_{A}(\gamma(t))$. Since $G^{\mathrm{F}}$ is right-invariant, $\gamma^{\prime}$ is also a null geodesic. Projecting $\gamma^{\prime}$ to $S\Sigma$, we obtain a chain ($g(\eta_{0},\dot{x}(a))\neq 0$) or a null chain ($g(\eta_{0},\dot{x}(a))=0$) through $\eta_{0}$. The uniqueness also follows from that of $\tilde{x}(t)$ and the fact that a lift of chain to a null geodesic is uniquely determined by the initial value. ∎
5.2. The Kropina metric for chains on $S\Sigma$ and a variational characterization of conformal geodesics
Hereafter, we identify the twistor CR manifold $M$ with the unit tangent sphere bundle $S\Sigma$ for $g\in[g]$, and consider the Fefferman space $(\mathcal{P},G^{\mathrm{F}})$ instead of $(\widehat{\mathcal{C}},g^{\mathrm{F}})$. We also write $K$ for the infinitesimal generator of the $S^{1}$-action on $\mathcal{P}$.
@@ 5.2. The Kropina metric for chains on $S\Sigma$ and a variational characterization of conformal geodesics
$F(\tilde{x},\xi):=\frac{G^{\mathrm{F}}(s_{*}\xi,s_{*}\xi)}{G^{\mathrm{F}}(K,s_{*}\xi)}\quad(\tilde{x}\in U,\ \xi\in T_{\tilde{x}}S\Sigma\setminus\mathrm{Ker}\,\theta)$
@@
We note that the condition $\xi\notin\mathrm{Ker}\,\theta$ is equivalent to $G^{\mathrm{F}}(K,s_{*}\xi)\neq 0$, and $F$ is not defined on the contact distribution.
@@
A curve $\tilde{x}(t)$ on $U\subset S\Sigma$ which is transverse to the contact distribution is a chain if and only if it is an (unparametrized) geodesic of $F$.
For a regular curve $x(t)\ (a\leq t\leq b)$ on $\Sigma$, we consider the Kropina length of the lift $\tilde{x}(t):=(x(t),\dot{x}(t)/|\dot{x}(t)|)$ associated with a section $s\colon U\to\mathcal{P}$ on a neighborhood $U\subset S\Sigma$ of the curve $\tilde{x}(t)$, and define the functional
@@
Noting that
$\phi(s_{*}\dot{\tilde{x}})=\begin{pmatrix}0\\ 0\\ |\dot{x}|\end{pmatrix},\quad\phi(K)=0,\quad\omega(K)=Y:=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&0\end{pmatrix},$
we have
$F(\tilde{x}(t),\dot{\tilde{x}}(t))=2\omega^{3}(s_{*}\dot{\tilde{x}}).$
If $x(t)$ is a conformal geodesic, $\tilde{x}(t)$ is a chain by Theorem 5.4, and hence $x(t)$ is a critical curve of $\mathscr{F}$ by Theorem 5.6. To prove the converse, we will show that the variation of the Kropina length in the vertical directions always vanishes:
@@
Let $\gamma(t)=(x(t),B(t))\ (a\leq t\leq b)$ be a curve on $\mathcal{P}$ which projects to $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$. Let $\gamma_{\epsilon}(t)$ be a variation of $\gamma(t)$ fixing the endpoints such that the variation vector field $(\partial/\partial\epsilon)|_{\epsilon=0}\gamma_{\epsilon}(t)$ is vertical with respect to the fiberation $\mathcal{P}\to\Sigma$. Then, we have
[…]
@@
(5.3) $\phi(\dot{\gamma}(t))=\begin{pmatrix}0\\ 0\\ |\dot{x}|\end{pmatrix}.$
@@
$\gamma_{\epsilon}(t)=R_{e^{\epsilon X(t)}}\gamma(t)=\gamma(t)\cdot e^{\epsilon X(t)},$
@@
$X(t)=\begin{pmatrix}0&X^{3}(t)&-X^{2}(t)\\ -X^{3}(t)&0&X^{1}(t)\\ X^{2}(t)&-X^{1}(t)&0\end{pmatrix}$
@@
$X^{\dagger}_{u}:=\frac{\partial}{\partial s}\Big{|}_{s=0}R_{e^{sX}}u,\quad u\in\mathcal{P}.$
@@
$(R_{A})_{*}X^{\dagger}=({\rm Ad}(A^{-1})X)^{\dagger}$
@@
$\dot{\gamma}_{\epsilon}(t)=(R_{e^{\epsilon X(t)}})_{*}\dot{\gamma}(t)+\epsilon\dot{X}(t)^{\dagger}.$

### v4

@@ 4.6 Null geodesics of $\boldsymbol{g^{\mathrm{F}}}$
Since $\widehat{f}$ maps any (null) chain on $M_{1}$ to a (null) chain on $M_{2}$, $f$ maps conformal geodesics on $\Sigma_{1}$ to conformal geodesics on $\Sigma_{2}$. Hence, by [20, Theorem 2.3], $f$ is a conformal isomorphism. ∎
4.7 Canonical lift of conformal geodesics
Theorem 1.1 (2) asserts that we can lift a conformal geodesic $x(t)$ to a unique (null) chain on $M$ once we choose a lift of the endpoint $x(a)$; we prove this in Section 5.1 below. Here, as a special case, we present a canonical lift of a conformal geodesic to a chain.
Given a regular curve $x(t)$ on $\Sigma$, we take a complex vector field $\zeta(t)$ along $x(t)$ such that $(\sqrt{2}\operatorname{Re}\zeta,\sqrt{2}\operatorname{Im}\zeta)$ gives an oriented orthonormal basis of $\dot{x}^{\perp}$ with respect to a metric $g\in[g]$}. Such a $\zeta(t)$ is determined by the conformal structure $[g]$ up to $\mathbb{C}^{*}$-action, so we have a canonical lift
@@ 4.7 Canonical lift of conformal geodesics
The curve $\tilde{x}(t)$ is a chain if and only if $x(t)$ is a conformal geodesic. Moreover, for a conformal geodesic $x(t)$, $a\leq t\leq b$, the curve $\tilde{x}(t)$ is a unique chain through $(x(a),[\zeta(a)])$ which projects to $x(t)$.
@@
$\eta=\dot{x},\qquad\dot{\gamma}^{0}=1,\qquad\dot{\gamma}^{1}=\dot{\gamma}^{2}=\dot{\gamma}^{3}=0.$
@@
Next, we prove the uniqueness. Suppose that $c(t)$ is a chain through $(x(a),[\zeta(a)])$ which projects to a conformal geodesic $x(t)$, and let $\gamma_{c}(t)=(x(t),\zeta_{c}(t))$ be a null geodesic on $\bigl(\widehat{\mathcal{C}},g^{\mathrm{F}}\bigr)$}, the Fefferman space determined by $g$ above, which projects to $c(t)$. Then, $\dot{\gamma}_{c}^{0}=g(\eta_{c},\dot{x})$ and $\dot{\gamma}_{c}^{1}=g(\zeta_{c},\dot{x})$ are constant by the geodesic equation. Since $\dot{\gamma}_{c}^{0}(a)=|\dot{x}(a)|^{2}=1$, $\dot{\gamma}_{c}^{1}(a)=0$, we obtain that $\eta_{c}(t)=\dot{x}(t)$ for any $t$ and hence $c(t)=\tilde{x}(t)$. ∎
5 A variational principle for conformal geodesics
in dimension three
5.1 Relations to the unit tangent sphere bundle and the frame bundle
We will give an alternative description of the previous constructions in terms of the frame bundle over $\Sigma$. We fix a representative metric $g\in[g]$ and let $\mathcal{P}$ be the $\mathrm{SO}(3)$-bundle of oriented orthonormal frames over $(\Sigma,g)$. Then, the map
$f\colon\ \widehat{\mathcal{C}}\overset{\sim}{\longrightarrow}\mathcal{P},\qquad\zeta\longmapsto B=\bigl(\sqrt{2}\operatorname{Re}\zeta,\sqrt{2}\operatorname{Im}\zeta,\eta\bigr)$ (5.1)
@@ 5.1 Relations to the unit tangent sphere bundle and the frame bundle
$M\overset{\sim}{\longrightarrow}S\Sigma,\qquad[\zeta]\longmapsto\eta.$ (5.2)
@@
$\phi=\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix}\colon\ T\mathcal{P}\longrightarrow\mathbb{R}^{3},\qquad\omega=\begin{pmatrix}\hphantom{-}0&\hphantom{-}\omega^{3}&-\omega^{2}\\ -\omega^{3}&\hphantom{-}0&\hphantom{-}\omega^{1}\\ \hphantom{-}\omega^{2}&-\omega^{1}&\hphantom{-}0\end{pmatrix}\colon\ T\mathcal{P}\longrightarrow\mathfrak{so}(3)$
be the canonical $1$-form and the Levi-Civita connection form of $g$. These forms are related to the adapted coframe $\bigl\{\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\bigr\}$ on $\widehat{\mathcal{C}}$ as follows.
@@
$f^{*}\phi=\begin{pmatrix}\sqrt{2}\operatorname{Re}\tilde{\theta}^{1}\\ \sqrt{2}\operatorname{Im}\tilde{\theta}^{1}\\ \theta\end{pmatrix},\qquad f^{*}\omega=\begin{pmatrix}0&\tilde{\theta}^{3}&-\sqrt{2}\operatorname{Im}\tilde{\theta}^{2}\\ -\tilde{\theta}^{3}&0&\hphantom{-}\sqrt{2}\operatorname{Re}\tilde{\theta}^{2}\\ \sqrt{2}\operatorname{Im}\tilde{\theta}^{2}&-\sqrt{2}\operatorname{Re}\tilde{\theta}^{2}&0\end{pmatrix}$.
@@
$f^{*}\phi={}^{t}\!B\,{\rm d}x=\begin{pmatrix}\sqrt{2}\operatorname{Re}{}^{t}\zeta\\ \sqrt{2}\operatorname{Im}{}^{t}\zeta\\ {}^{t}\eta\end{pmatrix}\begin{pmatrix}{\rm d}x^{1}\\ {\rm d}x^{2}\\ {\rm d}x^{3}\end{pmatrix}$
@@
$f^{*}\omega={}^{t}\!B\,{\rm d}B=\begin{pmatrix}\sqrt{2}\operatorname{Re}{}^{t}\zeta\\ \sqrt{2}\operatorname{Im}{}^{t}\zeta\\ {}^{t}\eta\end{pmatrix}\begin{pmatrix}\sqrt{2}\operatorname{Re}{\rm d}\zeta&\sqrt{2}\operatorname{Im}{\rm d}\zeta&{\rm d}\eta\end{pmatrix}$,
@@
We define the Fefferman metric $G^{\mathrm{F}}$ on $\mathcal{P}$ via the isomorphism (5.1): $g^{\mathrm{F}}=f^{*}G^{\mathrm{F}}$. Then, the above lemma and Theorem 4.5 prove the following.
@@
$G^{\mathrm{F}}=2\bigl(\phi^{1}\cdot\omega^{1}+\phi^{2}\cdot\omega^{2}+\phi^{3}\cdot\omega^{3}\bigr).$
@@
  • (i)
    If we denote by $R_{A}$ the right action of $A\in\mathrm{SO}(3)$ on $\mathcal{P}$, we have
    $R_{A}^{*}\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix}=A^{-1}\begin{pmatrix}\phi^{1}\\ \phi^{2}\\ \phi^{3}\end{pmatrix},\qquad R_{A}^{*}\begin{pmatrix}\omega^{1}\\ \omega^{2}\\ \omega^{3}\end{pmatrix}=A^{-1}\begin{pmatrix}\omega^{1}\\ \omega^{2}\\ \omega^{3}\end{pmatrix}$.
@@
  • (ii)
    Since the conformal class of the Fefferman metric is invariantly defined from the CR structure, the conformal manifold $\bigl(\mathcal{P},\bigl[G^{\mathrm{F}}\bigr]\bigr)$ should also be canonically associated to $(\Sigma,[g])$. This can be checked directly as follows. Let $\widehat{g}={\rm e}^{2\Upsilon}g$ be another representative metric and $\widehat{\mathcal{P}}$ be the oriented orthonormal frame bundle for $\widehat{g}$. Then, we have the canonical bundle isomorphism
    $\delta\colon\ \mathcal{P}\overset{\sim}{\longrightarrow}\widehat{\mathcal{P}},\qquad B\longmapsto{\rm e}^{-\Upsilon}B.$
    We denote by $\widehat{\phi}$, $\widehat{\omega}$ the canonical 1-form and the connection form of $\widehat{g}$ on $\widehat{\mathcal{P}}$. Then, we have $\delta^{*}\widehat{\phi}={\rm e}^{\Upsilon}\phi$ and the conformal transformation formula for the Levi-Civita connection gives
@@
    $\delta^{*}\widehat{G}^{\mathrm{F}}={\rm e}^{\Upsilon}G^{\mathrm{F}}.$
  • (iii)
    The conformal metric $G^{\mathrm{F}}$ on $\mathcal{P}$ is also considered by Holland [16], and it is proved in [16, Section 3.5.4] that null geodesics of this metric project to conformal geodesics on $\Sigma$. This gives another proof to Theorem 4.10.
In Section 4.7, we defined a canonical lift of a regular curve $x(t)$ on $\Sigma$ to a curve $\tilde{x}(t)=(x(t),[\zeta(t)])$ on $M$ which becomes a chain if and only if $x(t)$ is a conformal geodesic. This lift corresponds to the curve $(x(t),\dot{x}(t)/|\dot{x}(t)|)$ on $S\Sigma$ via the isomorphism (5.2). Thus, if we regard $S\Sigma$ as the twistor CR manifold, Theorem 4.12 yields the following.
@@
For a regular curve $x(t)$ on $(\Sigma,g)$, the curve $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$ on $S\Sigma$ is a chain if and only if $x(t)$ is a conformal geodesic of $[g]$. Moreover, for a conformal geodesic $x(t)$, $a\leq t\leq b$, the curve $\tilde{x}(t)$ is a unique chain through $(x(a),\dot{x}(a)/|\dot{x}(a)|)$ which projects to $x(t)$.
We can compete the proof of Theorem 1.1 (2) by making use of the right $\mathrm{SO}(3)$-invariance of $G^{\mathrm{F}}$.
@@
Let $x(t)$, $a\leq t\leq b$, be a conformal geodesic on $\Sigma$ and $\eta_{0}\in S_{x(a)}\Sigma$. Take a null geodesic $\gamma(t)$ on $\bigl(\mathcal{P},G^{\mathrm{F}}\bigr)$ which projects to the chain $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$. We choose $A\in\mathrm{SO}(3)$ so that the third basis vector in $R_{A}(\gamma(a))$ coincides with $\eta_{0}$, and set $\gamma^{\prime}(t):=R_{A}(\gamma(t))$. Since $G^{\mathrm{F}}$ is right-invariant, $\gamma^{\prime}$ is also a null geodesic. Projecting $\gamma^{\prime}$ to $S\Sigma$, we obtain a chain ($g(\eta_{0},\dot{x}(a))\neq 0$) or a null chain ($g(\eta_{0},\dot{x}(a))=0$) through $\eta_{0}$. The uniqueness also follows from that of $\tilde{x}(t)$ and the fact that a lift of chain to a null geodesic is uniquely determined by the initial value. ∎
5.2 The Kropina metric for chains on $\boldsymbol{S\Sigma}$ and a variational
characterization of conformal geodesics
Hereafter, we identify the twistor CR manifold $M$ with the unit tangent sphere bundle $S\Sigma$ for $g\in[g]$, and consider the Fefferman space $\bigl(\mathcal{P},G^{\mathrm{F}}\bigr)$ instead of $\bigl(\widehat{\mathcal{C}},g^{\mathrm{F}}\bigr)$. We also write $K$ for the infinitesimal generator of the $S^{1}$-action on $\mathcal{P}$.
@@ 5.2 The Kropina metric for chains on $\boldsymbol{S\Sigma}$ and a variational
$F(\tilde{x},\xi):=\frac{G^{\mathrm{F}}(s_{*}\xi,s_{*}\xi)}{G^{\mathrm{F}}(K,s_{*}\xi)},\qquad\tilde{x}\in U,\quad\xi\in T_{\tilde{x}}S\Sigma\setminus\operatorname{Ker}\theta$
@@
We note that the condition $\xi\notin\operatorname{Ker}\theta$ is equivalent to $G^{\mathrm{F}}(K,s_{*}\xi)\neq 0$, and $F$ is not defined on the contact distribution.
@@
A curve $\tilde{x}(t)$ on $U\subset S\Sigma$ which is transverse to the contact distribution is a chain if and only if it is an ((unparametrized)) geodesic of $F$.
For a regular curve $x(t)$, $a\leq t\leq b$, on $\Sigma$, we consider the Kropina length of the lift $\tilde{x}(t):=(x(t),\dot{x}(t)/|\dot{x}(t)|)$ associated with a section $s\colon U\to\mathcal{P}$ on a neighborhood $U\subset S\Sigma$ of the curve $\tilde{x}(t)$, and define the functional
$\mathscr{L}[x(t)]:=\int_{a}^{b}F\bigl(\tilde{x}(t),\dot{\tilde{x}}(t)\bigr)\,{\rm d}t.$
Noting that
$\phi(s_{*}\dot{\tilde{x}})=\begin{pmatrix}0\\ 0\\ |\dot{x}|\end{pmatrix},\qquad\phi(K)=0,\qquad\omega(K)=Y:=\begin{pmatrix}\hphantom{-}0&1&0\\ -1&0&0\\ \hphantom{-}0&0&0\end{pmatrix}$,
we have
@@
If $x(t)$ is a conformal geodesic, $\tilde{x}(t)$ is a chain by Theorem 5.4, and hence $x(t)$ is a critical curve of $\mathscr{L}$ by Theorem 5.6. To prove the converse, we will show that the variation of the Kropina length in the vertical directions always vanishes.
@@
Let $\gamma(t)=(x(t),B(t))$, $a\leq t\leq b$, be a curve on $\mathcal{P}$ which projects to $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$. Let $\gamma_{\epsilon}(t)$ be a variation of $\gamma(t)$ fixing the endpoints such that the variation vector field $(\partial/\partial\epsilon)|_{\epsilon=0}\gamma_{\epsilon}(t)$ is vertical with respect to the fiberation $\mathcal{P}\to\Sigma$. Then, we have
$\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Big|_{\epsilon=0}\int_{a}^{b}\frac{G^{\mathrm{F}}(\dot{\gamma}_{\epsilon}(t),\dot{\gamma}_{\epsilon}(t))}{G^{\mathrm{F}}(K,\dot{\gamma}_{\epsilon}(t))}\,{\rm d}t=0.$
@@
$\phi(\dot{\gamma}(t))=\begin{pmatrix}0\\ 0\\ |\dot{x}|\end{pmatrix}$. (5.3)
@@
$\gamma_{\epsilon}(t)=R_{{\rm e}^{\epsilon X(t)}}\gamma(t)=\gamma(t)\cdot{\rm e}^{\epsilon X(t)},$
@@
$X(t)=\begin{pmatrix}\hphantom{-}0&\hphantom{-}X^{3}(t)&-X^{2}(t)\\ -X^{3}(t)&\hphantom{-}0&\hphantom{-}X^{1}(t)\\ \hphantom{-}X^{2}(t)&-X^{1}(t)&\hphantom{-}0\end{pmatrix}$
@@
$X^{\dagger}_{u}:=\frac{\partial}{\partial s}\biggl|_{s=0}R_{{\rm e}^{sX}}u,\qquad u\in\mathcal{P}.$
@@
$(R_{A})_{*}X^{\dagger}=\bigl({\rm Ad}\bigl(A^{-1}\bigr)X\bigr)^{\dagger}$
@@
$\dot{\gamma}_{\epsilon}(t)=(R_{{\rm e}^{\epsilon X(t)}})_{*}\dot{\gamma}(t)+\epsilon\dot{X}(t)^{\dagger}.$